# A machine-learning photometric classifier for massive stars in nearby galaxies

### v1

@@ B.4 IC 10
The number of spectroscopically confirmed classified sources for IC 10 was 137 derived from five different papers. Of these sources we did not find a good match for 79 sources (57.7%). We had 11 sources passing the quality criteria resulting in 63.6% (7 sources) predicted correctly and 36.4% (4 sources) misclassified. By considering all (58) sources we got 39.7% (23 sources) as exact matches, 6.9% (4) as matches within errors, 50.0% (29) as erroneous matches, and 3.4% (2) as unclassified. By combining the correct matches we have a total success rate of 46.6%.
B.5 NGC 247
@@
There are 64 sources in NGC 247 with secured spectra derived from three works. Out of these, 17 sources ($\sim 26.6$%) do not have a good match within 1″. There are 11 sources that fulfilled the selection criteria and we successfully predicted their class for 9 of these (81.8%), while the other two sources were erroneously classified (18.2%). The total success rate is 81.8%. By considering all (47) sources we got 44.7% (21 sources) as exact matches, 14.9% (7) as matches within errors, 27.7% (13) as erroneous matches, and additionally 12.8% (6) uncertain objects (classified originally as ”star”, ”emission objects” without any further indication of any potential stellar signal, and composite). By combining the correct matches we have a total success rate of 59.6%.
B.6 NGC 253
We were able to gather 81 classified sources from four different works. 18 sources (22.2%) were removed because of no matching with any source from our catalog. Of the remaining 63 sources only 6 satisfied the quality criteria. From these we correctly predicted 16.7% (1 source), while 50.0% (3) were erroneously classified. There were additionally 33.3% (2 sources) that were uncertain. By considering the whole sample, we got 19.0% (12 sources) as exact matches, 15.9% (10) as matches within errors, 30.2% (19) as erroneous matches, and 34.9% (22) as uncertain. The total correct matches equals 34.9%. The rather low success rate is both due to the larger number of incorrect classifications and the fact that there are many more sources in this galaxy for which their classifications are not well constrained or they refer to objects for which the classifier has not been trained (e.g. ’cluster’ or ’nebula’).
B.7 NGC 300
@@
A total of 793 sources were collected from the literature. However, we have to remove the vast majority of them (553 sources, i.e. 69.7%) because they did not have any good match within 1″ of our catalog. Due to the lack of Pan-STARRS coverage there are no sources that can satisfy the band completeness criterion (¿0.6). Considering all (240) sources we got 10.4% (25) as exact matches, and 5.4% (13) as matches within errors, summing to 15.8%. A fraction of 80.8% (194) was predicted erroneously, while 3.3% (8) were uncertain cases. Due to the many missing features in this galaxy, the fraction of mis-prediction is high.
B.8 NGC 1313
About 100 classified sources were collected from two separate works in the literature. The majority of them (79.4%, 77 sources) lack a good match with our catalog, as well as properties above the selection criteria we set. Therefore, by considering all sources (20) we got exact matches for 35.0% (7 sources), matches within the error for 40.0% (8), erroneous predictions for 15.0% (3), and 10.0% (2) which we could not classify (uncertain). Overall, the total correct fraction is 75.0%.
B.9 IC 1613
@@
We collected 86 sources from 7 different works. About 64% of these (55 sources) do not have a good match with any of our catalog objects. Of the remaining 32 objects, 23 are good candidates (according to the selection criteria) and we managed to get 78.3% (18 sources) as exact matches, and 21.7% (5) erroneously classified. By considering all sources we got 64.5% (20 sources) as exact matches, 3.2% (1) as a good match within the errors, and 32.3% (10) as misclassifications. The total fraction of correctly predicted sources is 67.7%.
B.10 M33
@@
For this galaxy we have collected 1547 sources from 22 different works. About 61.3% (948) did not have any good match with our catalog. Of the remaining 599 sources, the majority (435) is of sufficient quality, and in particular 92.9% (404) were exact matches, and 7.1% (31) were erroneous. Considering all sources (599) we got 81.8% (490 sources) as exact matches, 6.2% (37) as a good match within the errors, 11.9% (71) as misclassifications, and 0.2% (1) as uncertain. The total fraction of correctly predicted sources is 88.0%.
B.11 Phoenix Dwarf
@@
Only eight objects with spectroscopic classifications were obtained from three different works in the literature. Unfortunately, the majority of them (75.0%, i.e. 6 sources) do not have any match within the 1″  search radius. Moreover, these two remaining sources do not pass the quality criteria and they are not predicted correctly. Therefore, we got a total correct fraction of 0%, which is due to the small number statistics and the fact that the band completeness (in all of these cases) is very low (at 0.2) which means that most of the features are missing.
B.12 NGC 1313
@@
We collected 97 sources from two different works. The majority of them (79.4%, 77 sources) do not have any good match with out catalog. Of the remaining 20 sources there are no good candidates (fulfilling the selection criteria). We had 35.0% (7 sources) as exact matches, 40.0% (8) as matches within the errors, 15.0% as misclassifications, and 10.0% (2) as uncertain. Therefore, the total correct fraction of sources was 75.0%.
B.13 NGC 2403
@@
We identified 68 sources originating from two works (Humphreys et al. 2019; Bresolin et al. 2022), and an additional (tentative) case (Kaldybekova et al. 2023). Unfortunately, 53 ($\sim 78\%$) of these sources (including the tentative case) do not have any good match with our catalog. From the remaining 15 sources 7 were of sufficient good quality and we got 28.6% (2) as exact matches, and 71.4% (5) as erroneous. When we considered all sources we got 20.0% (3) as exact matches, 26.7% (4) as matches within the errors, and 53.3% (8) as misclassifications. The total correct fraction of sources in this case was 46.7%.
B.14 M81
@@
A total of 82 sources were collected from the four works in the literature (including a number of AGN/galaxies and Hii regions). A good match was not found for the 67.1% (55) sources. From the remaining 27 sources, 12 passed the quality criteria, and we got 66.7% (8) as exact matches, and 33.3% (4) as erroneous. When considering all 27 sources we got 55.6% (15) as exact matches, 7.4% (2) as matches within the errors [^19] We note here that the catalog of Gómez-González et al. (2020) includes two separate sources, namely WR-3 and WR-14, which are reported as different sources with a separation of 0.87”. We opted not to remove any of them from our comparison list, although our search radius is 1″  . Therefore, these two sources are matched with the same object from our catalog (with ID M81-471) and it is a duplicate result. The impact is not significant though since this means a 3.6% decrease in the correct within the error and the total correct fraction., and 37.0% (10) as misclassifications. The total correct fraction in this case was 63.0%.
B.15 Sextans B
@@
Spectroscopic classifications hardly exist for this galaxy. There are a few objects studied by Britavskiy et al. (2019), but the majority are foreground sources (and therefore excluded from further consideration). We were left with only two sources, of which one had a match within 1″ . This source fulfilled the section criteria and it was actually predicted correctly (leading to an obviously biased fraction of 100%).
B.16 NGC 3109
@@
A total of 127 sources were gathered from five different works in the literature (including a number of Hii regions and planetary nebulae). A good match was not found for the majority of the sources (74.9%, 94 sources). From the remaining 33 sources, 15 passed the quality criteria, and we got 66.7% (10) as exact matches, 26.7% (4) as erroneous, and 6.7% (1) as uncertain. When considering all available sources we got 60.6% (20) as exact matches, 6.1% (2) as matches within the errors, 30.3% (10) as erroneous, and 3.0% (1) uncertain. The total correct fraction in this case was 66.7%.
B.17 Sextans A
@@
We collected 143 classified sources from 7 different works. A significant fraction of 63.6% of them (91 sources) were found more than 1″  away from our catalog sources. There were 14 sources with quality data of which we got 64.3% (9) as exact matches, and 35.7% (5) as erroneous. By considering all 52 sources, we got 28.8% (15) as exact matches, 30.8% (16) matches within the error, and 40.4% (21) as erroneous. The total success rate is 59.6%.
B.18 NGC 4736
@@
We only found three sources from the literature. Interestingly all of them are candidate LBVs (Solovyeva et al. 2019) of which one has been confirmed based on photometric and spectroscopic variability (Solovyeva et al. 2021). Unfortunately none of them has a good match with any of our catalog sources.
B.19 M83
We managed to obtain 241 sources from four different works. Unfortunately, the majority of them (185, 76.8%) are not matched with our catalog sources. Out of the remaining 56 objects only 8 fulfilled the selection criteria. We got 62.5% (5 sources) as exact matches, 25.0% (2) as erroneous and 12.5% (1) as uncertain. Considering all (56) objects, we got 57.1% (32) as exact matches, 14.3% (8) as correctly predicted within the error, 16.1% (9) as erroneous predictions, and 12.5% (7) as uncertain (due to the classification as clusters). The total correct fraction of our prediction was 71.4%.
B.20 NGC 6822
@@
We obtained 83 sources from the literature (6 different works). About 31% (26 sources) did not have a good match. Of the remaining 57 sources, 31 passed the quality criteria. We actually predicted all of them (96.8%, 30 source) correctly, without any misclassification, but only 1 source (3.2%) remained as unclassified. This can be justified because the spectral types of these sources are RSG and we have the best success rate for this particular class. Considering all sources, we got 57.9% (33 sources) correctly classified, 22.8% (13) correct within the classification error, 15.8% (9) misclassified, and another 3.5% (2) as uncertain. The total correct fraction in this case was 80.7%.
B.21 Pegasus DIG
@@
We found 9 sources from three different works. Two thirds of this sample (66.7%, 6 sources) did not have any match within 1″ . From the remaining 3 objects none fulfilled the selection criteria. Of these we got one predicted accurately, one correct within the classification error, and an erroneous one. Therefore, the success rate was at 66.7%, but it is based on a very small number of sources.
B.22 NGC 7793
@@
We collected 166 classified sources from 7 different works. A significant fraction (65.7%, 109 sources) did not have any good match with our catalog. From the remaining 57 sources only two had quality data, and both of them were erroneously predicted (leading to a 0% success rate). Considering all sources we get 10.5% (6) as exact matches, 15.8% (9) as matches within the error, 66.7% (38) predicted erroneously, and 7.0% (4) as uncertain. The total success rate in this case was 26.3%.
Appendix C A Baeysian approach to determine uncertainties in fractions
@@
In this Appendix, we describe the Bayesian approach used to determine the uncertainties in the fractions of the correct predictions. The likelihood is the probability of classifying correctly $k$ sources out of the $n$ of those in a certain class, given the success probability $p$ of the classifier. Consequently, it is a binomial distribution:
@@
where $P(p)$ is the prior probability of the success rate, and $P(k|n)$ is the marginal likelihood which can be treated as a normalization constant since it does not depend on $p$.
@@
[…]
where the lower, $l$, and upper, $u$, bounds are such that they contain the mode, and have equal probability density except for cases where the interval includes the extrema of $p$, 0 or 1, in which it can only expand towards one tail. In our implementation, the HPDI is calculated by starting from the mode, and expanding the interval at lower or higher values of $p$, depending on the direction at which the posterior density is higher and the interval can expand if it reached one of the extrema.
For the prior we explore the uniform distribution, $P(p)=1$, as well as the Beta distribution which is the conjugate distribution of the binomial, and is often employed as prior on probabilities. For the latter case, since we aim at using a prior with location and spread representing previous results, we constructed an algorithm that finds the parameters $\alpha$ and $\beta$ of the Beta distribution that result in the desired mode and variance. We do this by numerically solving for the parameters in the system of equations:
@@
The code for the Bayesian analysis, and the search for the unimodal Beta distribution, as well as the documentiation, are provided in the following open source repository: https://github.com/kkovlakas/gaussfree.

### v2

@@ C.3 IC 10
The number of spectroscopically confirmed classified sources for IC 10 was 137 derived from five different papers. Of these sources, we did not find a good match for 79 sources (57.7%), and we excluded another four sources with ambiguous classification. We had 11 sources passing the quality criteria, resulting in 63.6% (7 sources) with secure classifications, 18.2% (2 sources) as candidates, and 18.2% (2 sources) as uncertain, with a total success rate of 81.8%. By considering all (54) sources we got 37.0% (20) secure sources, 24.1% (13) candidates, and 38.0% (21) uncertain ones. By combining the secure classifications and candidate numbers, we get a total success rate of 61.6%.
C.4 M31
As the closest and largest galaxy the number of spectroscopically confirmed objects we managed to collect summed to 1170 sources (combining 13 different works). A bit more than half of them (57.4%, 671 sources) do not have a good match. The majority of the remaining sources (382) passed the quality criteria, with 93.7% (358) being secure predictions, about 1.0% being candidates (4), and only 5.2% being uncertain (20). Therefore, the total success rate was 94.8% When considering all 499 sources, the fraction of secure predictions was 73.3% (366), with additional 18.6% (93) candidates and 8.0% (40) uncertain cases. The total correct fraction is 92.0%. These high fractions are not unexpected, since the classifier has been trained on (most) of these sources.
C.5 NGC 247
@@
There are 64 sources in NGC 247 with secured spectra derived from three works. Of these, 17 sources ($\sim 26.6$%) do not have a good match within 1″, and another six sources have ambiguous classifications. There are 11 sources that fulfilled the selection criteria and we made secure predictions for nine of these (81.8%), while the other two sources were uncertain (18.2%). The total success rate is 81.8%. By considering all (41) sources we got secure classifications for 22.0% (9 sources), 36.6% (15) as candidates, and 41.5% (17) as uncertain. By combining the secure and candidate classification, the total success rate is 58.5%.
C.6 NGC 253
We were able to gather 81 classified sources from four different works. Eighteen sources (22.2%) were removed because of no match with any source from our catalog, and another 22 due to ambiguous classifications from the literature. Of the remaining 40 sources, only 4 satisfied the quality criteria. From these we obtained secure predictions for 16.7% (1 source), while 75.0% (3) were uncertain. The success rate is 25.0%. By considering the whole sample (40 sources), we again got 2.5% (1 source) as secure, 20.0% (8) as candidates, and 77.5% (31) as uncertain. The total success rate in this case is 22.5%. Similarly to NGC 55, this galaxy also suffers from significant missing values (band completeness equal to 0.2), which results in the low success rate we obtained.
C.7 NGC 300
@@
A total of 793 sources were collected from the literature. However, we had to remove the vast majority of them (553 sources, i.e., 69.7%) because they did not have any good match within 1″ of our catalog. We also excluded nine sources with ambiguous classification from the literature. Due to the lack of Pan-STARRS coverage, there are no sources that can satisfy the band completeness criterion (¿0.6). Considering all (231) sources, we got no secure classifications but only candidates (15.2%, i.e., 35 sources). A fraction of 84.8% (196) was predicted as uncertain. This leads to a success rate of 15.2% solely from the candidates. Due to the many missing features in this galaxy, the fraction of uncertain predictions is quite high.
C.8 IC 1613
@@
We collected 86 sources from seven different works. About 64% of these (55 sources) do not have a good match with any of our catalog objects. Of the remaining 31 objects, 23 are good candidates (according to the selection criteria) and we managed to get 78.3% (18 sources) secure predictions, 8.7% candidates, and 13.0% (3) as uncertain. The total success rate is 87.0%. By considering all sources, we got 58.1% (18) secure predictions, 19.4% (6) candidates, and 22.6% (7) uncertain classifications. The total success rate is 77.4%.
C.9 M33
@@
For this galaxy, we have collected 1547 sources from 22 different works. About 61.3% (948) did not have a good match with our catalog, and we excluded one ambiguous source. Of the remaining 598 sources, the majority (435) are of sufficient quality, and in particular 92.4% (402) were secure predictions, along with 1.4% (6) candidates, and 6.2% (27) uncertain cases. This sums to 93.8%. Considering all sources (598) we got 67.4% (43 sources) secure predictions, 23.6% (141) as candidates, and 9.0% (54) as uncertain. The total fraction of secure and candidate sources is 91.0%. Similarly to M31, we see high success rates since these sources lie in M33, which has been used for the training.
C.10 Phoenix Dwarf
@@
Only eight objects with spectroscopic classifications were obtained from three different works in the literature. Unfortunately, the majority of them (75.0%, i.e., six sources) do not have any match within the 1″  search radius. Moreover, these two remaining sources do not pass the quality criteria and they were predicted as uncertain. The success rate of 0% is actually due to the small number statistics and the fact that the band completeness (in all of these cases) is very low (at 0.2), which means that most of the features are missing.
C.11 NGC 1313
@@
We collected 97 sources from two different works. The majority of them (79.4%, 77 sources) have no good match with our catalog, while another two have ambiguous classifications. Of the remaining 18 sources, there are no good candidates (i.e. that fulfill the selection criteria). We did not manage to securely predict any of them, but the majority (72.2%, 13 sources) were predicted as candidates and 27.8% (five sources) as uncertain. Therefore, the total success rate was 72.2%.
C.12 NGC 2403
@@
We identified 68 sources originating from two works (Humphreys et al. 2019; Bresolin et al. 2022), and an additional (tentative) case (Kaldybekova et al. 2023). Unfortunately, 53 ($\sim 78\%$) of these sources (including the tentative case) have no good match with our catalog. Of the remaining 15 sources, seven were of sufficiently good quality and we got 28.6% (2) secure classifications, 42.9% (3) as candidates, and 28.6% (2) as uncertain. This results in a success rate of 71.4%. When we considered all sources, we got 20.0% (3) secure, 60.0% (9) candidates, and 20.0% (3) uncertain predictions. The total success rate was 80.0%.
C.13 M81
@@
A total of 82 sources were collected from the four works in the literature (including a number of AGN/galaxies and Hii regions). No good match was found for the 67.1% (55) sources. Of the remaining 27 sources, 12 passed the quality criteria, and we got 66.7% (8) as secure classifications and 33.3% (4) as uncertain, resulting in a success rate of 66.7%. When considering all 27 sources we got 55.6% (15) secure, 7.4% (2) candidate [^18] We note here that the catalog of Gómez-González et al. (2020) includes two separate sources, namely WR-3 and WR-14, which are reported as different sources with a separation of 0.87”. We opted not to remove any of them from our comparison list, although our search radius is 1″  . Therefore, these two sources are matched with the same object from our catalog (with ID M81-471) and it is a duplicate result. The impact is not significant however, since this means a 3.6% decrease in the candidates and the total success rate., and 37.0% (10) uncertain classifications. The total success rate in this case was 63.0%.
C.14 Sextans B
@@
Spectroscopic classifications hardly exist for this galaxy. There are a few objects studied by Britavskiy et al. (2019), but the majority are foreground sources (and therefore excluded from further consideration). We were left with only two sources, of which one had a match within 1″ . This source fulfilled the section criteria and was actually predicted correctly (leading to an obviously biased fraction of 100%).
C.15 NGC 3109
@@
A total of 127 sources were gathered from five different works in the literature (including a number of Hii regions and planetary nebulae). No good match was found for the majority of the sources (74.9%, 94 sources). We also excluded a single ambiguous classification. Of the remaining 32 sources, 14 passed the quality criteria, and we got 50.0% (7) secure, 7.1% (1) candidate, and 42.9% (6) uncertain classifications. This resulted in a success rate of 57.1%. When considering all available sources, we got 21.9% (7) secure, 37.5% (12) candidate, and 40.6% (13) uncertain classifications. The total success rate in this case was 59.4%.
C.16 Sextans A
@@
We collected 143 classified sources from seven different works. A significant fraction of 63.6% of them (91 sources) were found more than 1″  away from our catalog sources. There were 14 sources with quality data, of which we got 57.1% (8) secure, 21.4% (3) candidate, and 21.4% (3) uncertain classifications, accounting for a success rate of 78.6%. By considering all 52 sources, we got 15.4% (8) secure, 42.3% (22) candidate, and 42.3% (22) uncertain classifications. The total success rate was 57.7%.
C.17 NGC 4736
@@
We only found three sources from the literature. Interestingly, all of them are candidate LBVs (Solovyeva et al. 2019), of which one has been confirmed based on photometric and spectroscopic variability (Solovyeva et al. 2021). Unfortunately, none of them had a good match with any of our catalog sources.
C.18 M83
We managed to obtain 241 sources from four different works. Unfortunately, the majority of them (185, 76.8%) are not matched with our catalog sources. We also excluded another seven sources with ambiguous classification from the literature. Of the remaining 49 objects, only seven fulfilled the selection criteria. We got 28.6% (2 sources) secure and 71.4% (5) uncertain classifications (with a success rate of 28.6%). Considering all (49) objects, we got 22.4% (11) secure, 16.3% (8) candidate, and 61.2% (30) uncertain classifications. The total success rate was 38.8%. This rate is quite low, because the majority of these sources correspond to Hii regions, which are consistent with a number of stellar sources (compatible with LBV, BSG, WR, and B[e]SG). When implementing categorical cross entropy all these classes are taken into account (with their corresponding probabilities), resulting at the end in less secure predictions.
C.19 NGC 6822
@@
We obtained 83 sources from the literature (six different works). About 31% (26 sources) did not have a good match, while another two have ambiguous classifications from the literature. Of the remaining 55 sources, 30 passed the quality criteria. We actually predicted almost all of them (96.7%, 30 sources) securely, and only one source (3.2%) as uncertain, i.e., a success rate of 96.7%. This can be justified because the spectral types of these sources are RSG and we have the best success rate for this particular class. Considering all sources, we got 52.7% (29 sources) as secure classifications, 23.6% (13) as candidates, and 23.6% (13) as uncertain. The total success rate was 76.4%.
C.20 Pegasus DIG
@@
We found nine sources from three different works. Two thirds of this sample (66.7%, 6 sources) did not have any match within 1″ . Of the remaining three objects, none fulfilled the selection criteria. Of these we got none with a secure classification, one candidate (33.3%) and two uncertain (66.7%) classifications. Therefore, the success rate was 33.3%, but is based on a very small number of sources.
C.21 NGC 7793
@@
We collected 166 classified sources from seven different works. A significant fraction (65.7%, 109 sources) had no good match with our catalog, while an additional four sources have ambiguous classifications. From the remaining 53 sources, only two had quality data, and both were predicted as uncertain (leading to a 0% success rate). Considering all sources, we got only 1.9%, i.e., one source as secure classification, 30.2% (16) as candidates, and 67.9% (36) as uncertain. The total success rate in this case was 32.1%.
Appendix D A Bayesian approach to determine uncertainties in fractions
@@
In this Appendix we describe the Bayesian approach used to determine the uncertainties in the fractions of the correct predictions. The likelihood is the probability of correctly classifying $k$ sources from the $n$ of those in a certain class, given the success probability $p$ of the classifier. Consequently, it is a binomial distribution:
@@
where $P(p)$ is the prior probability of the success rate, and $P(k|n)$ is the marginal likelihood, which can be treated as a normalization constant since it does not depend on $p$.
@@
$\int_{l}^{u}P(p|k,n)dp$ $=0.68,$ (4)
where the lower, $l$, and upper, $u$, bounds are such that they contain the mode, and have equal probability density except for cases where the interval includes the extrema of $p$, 0 or 1, in which it can only expand towards one tail. In our implementation, the HPDI is calculated by starting from the mode, and expanding the interval at lower or higher values of $p$, depending on the direction in which the posterior density is higher and the interval can expand if it reaches one of the extrema.
For the prior, we explore the uniform distribution, $P(p)=1$, as well as the Beta distribution, which is the conjugate distribution of the binomial, and is often employed as prior on probabilities. For the latter case, since we aim to use a prior with location and spread representing previous results, we constructed an algorithm that finds the parameters $\alpha$ and $\beta$ of the Beta distribution that result in the desired mode and variance. We do this by numerically solving for the parameters in the system of equations:
@@
The code for the Bayesian analysis and the search for the unimodal Beta distribution, as well as the documentation, are provided in the following open source repository: https://github.com/kkovlakas/gaussfree.